# UniGenX: Unified Generation of Sequence and Structure with Autoregressive Diffusion

### v1

@@ 4.1.1 Crystal Structure Prediction
In this task, we performed finetuning and evaluations on three mainstream benchmarks. The MP-20 dataset is a subset of the Materials Project [51]. It encompasses all materials with a unit cell atomic count less than 20, and most properties can achieve DFT accuracy. Carbon-24 [52] contains 10k carbon materials, which share the same composition, but have different structures. Similar to the MP-20 dataset, it covers materials composed only of carbon atoms with a unit cell atomic count less than 24. The MPTS-52 dataset includes time-based splits for cross-validation and benchmarking generative models, allowing up to 52 atoms in the structures.
And the results will be compared with three baselines which are all previous state-of-the-art methods. CDVAE [53] is a VAE-based framework designed for generating pure crystals. It is modified in [38] to perform the CSP task by replacing its original normal prior for generation with a parametric prior conditioned on the encoding of the specified composition. DiffCSP [38] is based on diffusion model. It jointly generates lattice vectors and fractional coordinates to learn the structure distribution of stable crystals. The latest SOTA metheod is FlowMM [39], which uses Riemannian Flow Matching to create distributions over fractional atomic coordinates, unit cells, and atomic types, ensuring invariance to translations, rotations, and permutations inherent in crystals. While our model does not incorporate explicit inductive biases for equivariance or invariance, we demonstrate through our results that these properties can be effectively learned through data augmentation, preserving the scalability of the underlying transformer architecture and providing flexibility across diverse tasks and scientific domains with varying symmetry requirements.
@@ 4.2.1 Conformation Generation
For evaluation, given molecule $x$ with $N_{x}$ conformations in the test set, we generate $2N_{x}$ conformations, following [58]. Let $\mathbb{S}_{g}$ and $\mathbb{S}_{r}$ represent the sets of generated and ground truth conformations, respectively. Conformation deviation is quantified using GetBestRMS from the RDKit package, with root-mean-square deviation denoted as $\mathrm{RMSD}(R,\hat{R})$. The recall-based coverage (COV) and matching (MAT) scores are defined as follows:
@@
$\mathrm{MAT}(\mathbb{S}_{g},\mathbb{S}_{r})=\frac{1}{|\mathbb{S}_{r}|}\sum_{R\in\mathbb{S}_{r}}{min}_{\hat{R}\in\mathbb{S}_{g}}\mathrm{RMSD}(R,\hat{R})$ (11)
@@
$\mathrm{MAT}\text{-}\mathrm{P}(\mathbb{S}_{g},\mathbb{S}_{r})=\frac{1}{|\mathbb{S}_{g}|}\sum_{{\hat{R}\in\mathbb{S}_{g}}}{min}_{R\in\mathbb{S}_{r}}\mathrm{RMSD}(R,\hat{R})$ (13)
4.2.2 Property Prediction
We extended our evaluation to include molecular property prediction alongside conformation generation, a task that involves predicting properties from generated conformations [55]. Following [36], we randomly selected 30 molecular SMILES from the test set of the GEOM-QM9 dataset. For each molecule, we generated 50 distinct conformations. Using the Psi4 quantum chemistry software package [60], we calculated energy levels, HOMO, and LUMO characteristics for both generated and reference conformations. We then derived ensemble averages for energy ($\overline{E}$), minimum energy ($E_{\text{min}}$), average HOMO-LUMO gap ($\Delta\overline{\epsilon}$), and the gap extremities ($\Delta\epsilon_{\text{min}}$ and $\Delta\epsilon_{\text{max}}$) based on the conformational attributes of each molecule. Mean absolute error (MAE) was used to quantify property discrepancies between generated and reference conformations. Given our model’s superior performance in predicting these statistical quantities, depicted in Sec. 5.2, such as average energy and average HOMO-LUMO gap, we further compared the distribution of sampled results from UniGenX with the ground-truth conformations. To our knowledge, this is the first metric of distribution comparison performed for this task.
@@ 4.2.3 Conditional Generation
We also explored our model’s conditional generation capabilities. Instead of employing guidance in the diffusion model, UniGenX utilizes a straightforward approach: prepending the training data sequence with the condition property and its value, marked by a special token like $<$bobulk$>$ for the bulk property. This enables simple prompting at inference, where conditional generation is achieved by providing the property-value pair and the sequence generation special token. Data processing examples are detailed in Appendix C.5. In this section, using conditional molecule generation on the QM9 benchmark as an example, we demonstrate UniGenX’s strong performance and note its easy extensibility to different domains.
@@
The properties include $\alpha(Bohr^{3})$, the ability of a molecule to become polarized under an external electric field; energy gap $\Delta\varepsilon(meV)$, the energy difference between HOMO and LUMO; HOMO $\varepsilon_{H}(meV)$, the energy level of the highest occupied molecular orbital; LUMO $\varepsilon_{L}(meV)$, the lowest unoccupied molecular orbital energy; dipole moment $\mu(D)$, the separation of positive and negative charges within the molecule.
For evaluation, we employed two different sampling methods. The first, introduced in [61], models the distribution of conditions in the training set based on 15. It first samples the number of atoms, $N$ , and then samples the property values corresponding to $N$. The second method, proposed by [62], addresses a potential inductive bias in the original method from [61], where an implicit correlation between molecule size and properties may arise during evaluation. To mitigate this, they propose uniformly sampling property values directly between the minimum and maximum values, as described in 14.
We evaluated our model’s performance aligning both sampling methods. Notably, unlike prior approaches that trained separate models for each property and relied on extensive preprocessing, our model is trained as a unified, all-in-one model, directly utilizing raw data without requiring pre-generation of the number of atoms, N.
@@ 4.3 Unification training on Material and Molecule
To differentiate domains, we introduced special tokens for materials and molecules as indicators during training. We trained a 100M parameter model on a combined dataset of 5M material data and 1M QM9 molecule data, denoted as UniGenX(100M). We then fine-tuned this model on material datasets including MP-20, Carbon-24, and MPTS-52, as well as the 1M GEOM-QM9 molecule dataset.
4.4 Combing with Natural Language Model
Building upon our approach, which leverages an autoregressive sequence model backbone and joint training on symbolic and numerical data, UniGenX naturally lends itself to language capabilities. We explored this by loading a pretrained language model and performing instruction tuning on the MP-20 crystal structure prediction task using language prompts. The MP-20 dataset was reconstructed into an instruction tuning dataset, with the data format detailed in Appendix C.6.
We then fine-tuned the pretrained model, which had approximately 1B parameters from NLM [35], on this dataset, demonstrating its superior performance.
@@ 5 Results
We evaluated UniGenX on crystal structure prediction and de novo generation in the material domain, conformation generation and conditional generation in the molecular domain, unification and language capacity. Table S2 summarizes the computational resources used, including model size, GPU type, and total GPU days.
@@ 5.1 Material Results
Methods MR(%$\uparrow$ RMSD $\downarrow$ MR(%$\uparrow$ RMSD $\downarrow$ MR(%$\uparrow$ RMSD $\downarrow$
@@
Table 1: Crystal structure prediction results on material benchmarks. “-” indicates unreported results. Our model significantly outperforms the state-of-the-art, with increased performance on longer sequences.
Methods InSteps VS (%$\uparrow$ VC (%$\uparrow$ CR (%$\uparrow$ CP(%$\uparrow$ PW($\rho$) $\downarrow$ PW(Nel) $\downarrow$
@@
Table 2: De novo generation tasks benchmark. IS, VS, VC, CR, CP and PW represent Integration steps, Validity Structural, Validity Composition, Coverage Recall, Coverage Precision and Property Wdist.
Table 1 and Table 2 present the evaluation metrics for baseline methods and our model in Crystal Structure Prediction and de novo generation, respectively. As shown in Table 1, our model significantly outperforms the previous diffusion/flow matching state-of-the-art, FlowMM, achieving match rate increases of 10%, 28%, and 120% on MP-20, Carbon-24, and MPTS-52, respectively. In terms of RMSD, we observe improvements of 35%, 45%, and 62%. Compared to NatureLM [35], a recently published autoregressive model that surpassed FlowMM on the MP-20 and MPTS-52 benchmarks, our model also demonstrates superior performance, establishing a new state-of-the-art across MP-20, Carbon-24, and MPTS-52.
@@
Some examples of crystal structure prediction are shown in Figure 3, and Figure 4 displays some examples of de novo generation.

### v2

@@ 3.1.1 Crystal Structure Prediction
In this task, we performed finetuning and evaluations on three mainstream benchmarks. The MP-20 dataset is a subset of the Materials Project [13]. It encompasses all materials with a unit cell atomic count less than 20, and most properties can achieve DFT accuracy. Carbon-24 [14] contains 10k carbon materials, which share the same composition, but have different structures. Similar to the MP-20 dataset, it covers materials composed only of carbon atoms with a unit cell atomic count less than 24. The MPTS-52 dataset includes time-based splits for cross-validation and benchmarking generative models, allowing up to 52 atoms in the unit cells.
The performance is compared with three baselines in Table 1 which are all previous state-of-the-art methods. Notably, In contrast to them, our model does not incorporate explicit inductive biases for equivariance or invariance, we demonstrate through our results that these properties can be effectively learned through data augmentation, preserving the scalability of the underlying transformer architecture and providing flexibility across diverse tasks and scientific domains with varying symmetry requirements. CDVAE [15] is modified in [9] to perform the CSP task by replacing its original normal prior for generation with a parametric prior conditioned on the encoding of the specified composition.
Method Model Type
CDVAE[15](modified by [9]) Variational Autoencoder (VAE)
DiffCSP[9] Diffusion Model
FlowMM[10] Riemannian Flow Matching
Table 1: Comparison of generative models for crystal structure prediction.
@@ 3.1.2 De Novo Generation
3.1.3 Conditional Generation
As an autoregressive model, one of its key features is the ability to perform controllable conditional generation with ease. Therefore, we evaluate UniGenX on conditional generation tasks, where the goal is to generate stable crystal structures that satisfy specific target conditions. This setting reflects real-world design scenarios in which materials are sought with desired physical or chemical properties.
To enable conditional generation, we prepend a special token to the input sequence, which includes both the material composition and structure, to indicate the target property type. Sequentially, we embed the target property value using a multilayer perceptron (MLP) and append it to the sequence, in front of the material’s composition and structure. We finetune our pretrained model using datasets with property labels. We focus on scalar property conditioning, targeting single-valued material properties such as magnetic density and bulk modulus. For each property, we train a separate model using a small-to-moderate-sized labeled dataset.
During training, we apply the loss only on the material itself, without directly supervising the property value. To better handle the numerical range of the target properties, we apply a nonlinear transformation of the form $\mathrm{sign}(x)\cdot\log(|x|+1)$ to the property values before embedding.
To demonstrate our model’s ability to generate out-of-distribution (OOD) samples, following [17], we consider two scalar property targets:
  • Magnetic Density: We finetune on 600k structures labeled with magnetic density (in $\text{\AA }^{-3}$), and generate samples targeting values equal to 0.20 $\text{\AA }^{-3}$.
  • Bulk Modulus: We finetune on a small dataset of 5k samples, targeting high-stiffness materials with a bulk modulus equal to 400 GPa.
During inference, inspired by [18], we employed an earlier checkpoint of our own model—trained for fewer steps—to guide the generation process of the final model. This approach serves as a substitute for the standard classifier-free guidance (CFG), effectively enabling a form of “temperature reduction” during sampling. For the generated materials, property values were evaluated using the MatterSim [19].
3.1.4 Multi-conditional Generation
We further evaluated UniGenX under multi-conditional generation, where the target is to generate materials satisfying multiple properties simultaneously. This is a classic materials design problem involving trade-offs between multiple physical properties. The core significance of this task lies in the fact that these three properties are typically in conflict with one another, making simultaneous optimization highly challenging—providing an ideal benchmark to evaluate the capability of a material generation model. Ultimately, these materials can potentially provide implications in many high-value engineering applications such as aerospace materials, thermal energy storage and management materials, and protective materials etc.
In this setting, we prepend three special tokens to the input sequence, each indicating a specific property type: heat capacity at 300 K (J/g·K), Hill’s elastic modulus(GPa), and density (kg/m³), followed by their corresponding signed-log-transformed values.
Training and inference procedures remain consistent with the single-condition setup, including the use of classifier-free guidance during sampling.
The multi-labeled dataset is constructed from the Materials Project containing no more than 16 atoms, and employed the pre-trained MatterSim-v1.0.0-5M machine-learning interatomic potential via the ASE and Phonopy interfaces to compute, under the harmonic approximation, each crystal’s Hill’s elastic modulus, its thermal conductivity at 300 K, and its density—these quantities serving as the training labels. For each data item containing three distinct labels, we partition it into $C_{3}^{1}+C_{3}^{2}+C_{3}^{3}$ subsets, where each subset contains 1, 2, or 3 labels, respectively. These subsets are then used to fine-tune our model. This strategy is designed to enable UniGenX to flexibly generate structures under varying numbers of conditioning labels, thereby enhancing its versatility in both under- and fully-conditioned generation scenarios.
As a representative multi-conditional generation task, we aim to generate crystal structures that satisfy the following target properties: density of $2.5\pm 0.2kg/m^{3}$, Hill’s elastic modulus E>200E>200GPa, and specific heat capacity $C_{v}>1$J/g·K.
@@ 3.2.1 Conformation Generation
[…]
@@
[…]
@@
[…]
3.2.2 Property Evaluation
We extended our evaluation to include molecular property prediction alongside conformation generation, a task that involves predicting properties from generated conformations [21]. Following [20], we randomly selected 30 molecular SMILES from the test set of the GEOM-QM9 dataset. For each molecule, we generated 50 distinct conformations. Using the Psi4 quantum chemistry software package [26], we calculated energy levels, HOMO, and LUMO characteristics for both generated and reference conformations. We then derived ensemble averages for energy ($\overline{E}$), minimum energy ($E_{\text{min}}$), average HOMO-LUMO gap ($\Delta\overline{\epsilon}$), and the gap extremities ($\Delta\epsilon_{\text{min}}$ and $\Delta\epsilon_{\text{max}}$) based on the conformational attributes of each molecule. Mean absolute error (MAE) was used to quantify property discrepancies between generated and reference conformations. Given our model’s superior performance in predicting these statistical quantities, depicted in Sec. 4.2, such as average energy and average HOMO-LUMO gap, we further compared the distribution of sampled results from UniGenX with the ground-truth conformations.
@@ 3.2.3 Conditional Generation
We further explored our model’s conditional generation capabilities. Instead of employing guidance in the diffusion model, UniGenX utilizes a straightforward approach: prepending the training data sequence with the condition property and its value, marked by a special token like <<bobulk>> for the bulk property. This enables simple prompting at inference, where conditional generation is achieved by providing the property-value pair and the sequence generation special token. Data processing examples are detailed in Appendix C.5. In this section, using conditional molecule generation on the QM9 benchmark as an example, we demonstrate UniGenX’s strong performance and note its easy extensibility to different domains.
@@
The properties include volume $\alpha(Bohr^{3})$, the ability of a molecule to become polarized under an external electric field; HOMO $\varepsilon_{H}(meV)$, the energy level of the highest occupied molecular orbital; LUMO $\varepsilon_{L}(meV)$, the lowest unoccupied molecular orbital energy; energy gap $\Delta\varepsilon(meV)$, the energy difference between HOMO and LUMO; dipole moment $\mu(D)$, the separation of positive and negative charges within the molecule.
For evaluation, we employed two different sampling methods. The first, introduced in [27], models the distribution of conditions in the training set based on equation 6. It first samples the number of atoms, $N$ , and then samples the property values corresponding to $N$. The second method, proposed by [28], addresses a potential inductive bias in the original method from [27], where an implicit correlation between molecule size and properties may arise during evaluation. To mitigate this, they propose uniformly sampling property values directly between the minimum and maximum values, as described in equation 5.
We evaluated our model’s performance aligning both sampling methods. Notably, unlike prior approaches that trained separate models for each property and relied on extensive preprocessing, our model is trained as a unified, all-in-one model and not require pre-generation of the number of atoms, N.
@@
3.2.4 Multi-conditional Generation
In the domain of small molecule conditional generation, another crucial task is multi-condition generation. Since our model is a sequence model, it easily accommodates multiple condition properties by appending them to the beginning of the sequence for conditional generation.
We have observed that previous works typically focus on single-condition generation and train models using only one condition. Here, we introduce a multi-condition generation setting. Consistent with [27], we divide the QM9 dataset equally into two halves, with the first half used to train the classifier and the second half to train our model. The QM9 dataset contains six comparable properties, so we select $n$ properties (where $1\leq n\leq 6$) for splitting the data. This results in a total of $\binom{6}{1}+\binom{6}{2}+\binom{6}{3}+\binom{6}{4}+\binom{6}{5}+\binom{6}{6}=63$ data splits. We continue to use the training weights provided by [28] for our classifier. Similar to single-condition generation, we embed conditions into the sequence, requiring only one special token per condition for embedding. For multi-condition generation, we determine the order of conditions based on the sorted string representation of them.
During the evaluation phase, we randomly selected 1k values from any of the 63 condition combinations in the divided test set for generation, and subsequently assessed them using a classifier. We believe that calculating the Mean Absolute Error (MAE) for each property in multi-condition generation lacks significant meaning. The primary focus is for our generator to produce distributions that meet real-world requirements. Therefore, for different condition combinations, we plotted the distribution of both the actual data and the generated data.
3.3 Protein
We pre-trained a 400M-parameter model using approximately 78 million protein sequence–structure pairs from the AFDB. These proteins were selected based on sequence identity below 90% and pLDDT scores above 70. During training, sequences longer than 256 residues were randomly truncated, enabling the model to effectively learn the relationship between protein sequences and their 3D structures, and to support downstream tasks.
3.3.1 Emulating MD equilibrium distributions
We evaluated UniGenX’s ability to model long-timescale molecular dynamics (MD) equilibrium distributions using simulations of 12 fast-folding proteins generated by [29] on the special-purpose supercomputer Anton.
[…]
We then trained 12 “DESRES-finetuned models,” each fine-tuned on 11 fast folders using our base model pretrained on AFDB data, and tested on the remaining held-out fast folder. Across these 12 tasks, our model successfully recovered native and unfolded states that are structurally consistent with the expected free energy landscapes of the proteins.
For evaluation, we employed Time-lagged Independent Component Analysis (TICA) to identify low-dimensional projections associated with the system’s slowest dynamical processes. Using pairwise distance matrices of protein structures as input, TICA extracts directions corresponding to the most kinetically relevant degrees of freedom, which effectively serve as latent axes of the free energy landscape. We then projected the trajectories onto the first two TICA components to visualize the system in a 2D plane, where each point represents a protein conformation. High-density regions in this space correspond to low free energy (i.e., highly probable) states, while sparse regions indicate high free energy (i.e., less probable) states. This analysis enables us to assess whether UniGenX has successfully learned to reproduce the true long-timescale MD equilibrium distributions.
3.3.2 EC Number-guided Protein Generation
The Enzyme Commission (EC) number is a numerical classification scheme for enzymes based on the chemical reactions they catalyze. Each EC number consists of four hierarchical levels (e.g., 2.6.1.1), providing increasingly specific information about enzyme function. In our work, we use only the first three levels (e.g., 2.6.1) as generation conditions, which is sufficient to define the primary functional class of the target enzyme while maintaining generalization.
Our approach to EC number-guided protein generation follows the same design as the molecular and material property-conditional generation described earlier. Specifically, the EC number is converted into a sequence of tokens and special tokens, which are prepended to the input sequence. For example, an EC number 2.6.1.x is encoded as <ec1>2<ec2>6<ec3>1<prot> at the beginning of the input sequence and structure tokens.
For training data, we used all protein sequences in UniProt that have an assigned EC number and corresponding structural information available in the AlphaFold Database (AFDB), and with sequence lengths below 512 residues. This results in a dataset of approximately 20 million samples.
To evaluate the effectiveness of EC number conditioning, we selected three representative EC numbers and generated 400 protein sequences for each. The validity of the generated sequences was first checked using CLEAN [30] to ensure they conformed to the assigned EC number. To further assess the structural quality and functional relevance, we predicted the 3D structures of the generated sequences using ESMFold [31], performed sequence alignment against the UniProt database using BLAST, and conducted structural comparisons against AFDB and PDB using Foldseek [32]. Additionally, to evaluate the contribution of structure information during training, we trained a sequence-only version of the model as a baseline for comparison.
3.4 Protein-Ligand Docking
A central challenge in structural biology is the accurate prediction of small-molecule ligands that bind to a given protein pocket[33]. Most existing approaches rely on static, co-crystallized protein-ligand complexes, failing to account for the conformational flexibility that proteins often exhibit upon ligand binding. In this work, we formulate two related tasks that explicitly incorporate protein dynamics into the modeling process.
The first task involves predicting the bound ligand conformation given both the unbound (apo) and bound (holo) pocket structures. The second, more challenging task requires predicting both the holo pocket conformation and the bound ligand coordinates from the apo structure alone. Together, these tasks aim to model the dynamic nature of molecular recognition and capture induced-fit effects that are often neglected in traditional structure-based docking pipelines. To this end, we leverage the MISATO dataset[34], which comprises approximately 20,000 experimentally resolved protein-ligand complexes along with over 170 microseconds of molecular dynamics (MD) trajectories—offering a unique view into the conformational transitions involved in binding events.
To construct inputs for these tasks, we use AlphaFold2-predicted structures[35] as proxies for the unbound (apo) state of proteins. Because the MISATO dataset does not retain original chain identifiers and includes preprocessing steps such as hydrogen addition, we use the PDBbind dataset[36] as a reference. We first align each MISATO complex to its corresponding entry in PDBbind, then align the relevant chains to their AlphaFold2-predicted structures from the AlphaFold Protein Structure Database (AFDB) using UniProt identifiers. This two-step alignment allows us to establish residue-level correspondences between experimental and predicted structures. Binding pockets are subsequently extracted using the MISATO-provided script, which selects all residues within 10 Å of the ligand.
Following this preprocessing pipeline, we obtain 9,744 protein-ligand complexes, each with 100 MD snapshots. We adopt the CASF-2016 benchmark split[37] to partition the data into 9,587 training samples and 157 test samples. Each small molecule is represented using its canonical SMILES string after hydrogen removal, while protein pockets are represented at the all-atom level.
Each test sample is evaluated using $N_{f} = 2$00 generated conformational frames. The two prediction tasks are assessed using distinct evaluation protocols. In the ligand-only prediction task, where both apo and holo pocket structures are provided as input, performance is measured by computing the root-mean-square deviation (RMSD) between the predicted ligand coordinates and the ground truth ligand in the corresponding MD frame. In the joint prediction task, which requires generating both the holo pocket and the ligand from the apo structure alone, evaluation is conducted separately for the ligand, the pocket, and the full protein-ligand complex. For each of these components, RMSD is computed against all MD frames corresponding to the same PDB ID, and the minimum RMSD across frames is reported. This flexible evaluation accounts for the conformational ambiguity inherent in dynamic binding processes, recognizing that multiple structurally distinct solutions may be equally valid.
@@ 3.5 Unification training on Material and Molecule
To differentiate domains, we introduced special tokens for materials and molecules as indicators during training. We trained a 100M parameter model on a combined dataset of 5M material data and 1M QM9 molecule data, denoted as UniGenX(100M). We then finetuned this model on material datasets including MP-20, Carbon-24, and MPTS-52, as well as the 1M GEOM-QM9 molecule dataset.
3.6 Combing with Large Language Model
To facilitate more effective integration with the LLM, we performed two tasks: (1) fine-tuning the language model to endow it with the capability of predicting crystal structures; and (2) incorporating a gating module that enables the model to dynamically switch between generating textual and numerical outputs.
Building upon our approach, which leverages an autoregressive sequence model backbone and joint training on symbolic and numerical data, UniGenX naturally lends itself to language capabilities. We explored this by loading a pretrained language model and performing instruction tuning on the structure prediction task using language prompts. For example, The MP-20 dataset was reconstructed into an instruction tuning dataset, with the data format detailed in Appendix C.10.
To enable the model to seamlessly switch between generating textual and numerical outputs — such as producing a set of structural coordinates followed by explanatory text — it is crucial for the model to autonomously predict whether the feature vector should be routed through the lm head or the diffusion head prior to entering them, rather than relying on this information as a given.
For this purpose, we introduced a gating module shown in Figure 4 that performs a binary classification on the model’s final hidden state to determine which head it should pass through.
We then finetuned the pretrained model, which had approximately 1B parameters from NatureLM [38], on MP-20 and QM9, demonstrating its superior performance.
Figure 3: An example for instruction tunning for the Crystal Structure Prediction (CSP) task.
Figure 4: Diagram of the hidden state, gate, and heads.
@@ 4 Results
We evaluated UniGenX on crystal structure prediction and de novo generation in the material domain, conformation generation and conditional generation in the molecular domain, unification and language capacity.
@@ 4.1 Material Results
[…]
@@
Table 2: Crystal structure prediction results on material benchmarks. “-” indicates unreported results. Compared to the latest, our model achieves a 10% improvement in match rate on MP-20, and a 28% improvement on Carbon-24 and MPTS-52. Our model significantly outperforms the state-of-the-art, with increased performance on longer sequences(materials with larger unit cells).
[…]
@@
Table 3: De novo generation tasks benchmark. The abbreviations in the table refer to: Integration steps(InSteps), Validity Structural(VS), Validity Composition(VC), Coverage Recall(CR), Coverage Precision(CP) and Property Wdist(PW).
The evaluation metrics are summarized in Table 2 and Table 3 for baseline methods and our model in Crystal Structure Prediction and de novo generation, respectively. As shown in Table 2, our model significantly outperforms the previous diffusion/flow matching state-of-the-art, FlowMM, achieving match rate increases of 10%, 28%, and 120% on MP-20, Carbon-24, and MPTS-52, respectively. In terms of RMSD, we observe improvements of 35%, 45%, and 62%. Compared to NatureLM [38], a recently published autoregressive model that surpassed FlowMM on the MP-20 and MPTS-52 benchmarks, our model also demonstrates superior performance, establishing a new state-of-the-art across MP-20, Carbon-24, and MPTS-52.
@@
Some examples of crystal structure prediction results are shown in Figure 5, and Figure 6 displays some examples of de novo generation, in terms of both material formula and crystal structures.
Figure 7: Generated materials with target magnetic and mechanical properties. Panels (a) and (e) show the density distributions of the target properties for the stable samples generated by UniGenX, alongside the corresponding label distributions in the fine-tuning dataset for magnetic (a) and mechanical (e) properties. The target value specified for UniGenX is indicated by a black dashed line. Panels (b–d) and (f–h) visualize representative S.U.N. structures generated under magnetic (b–d) and mechanical (f–h) property conditions, each annotated with the reduced formula, space group, and the predicted property value. The model demonstrates the ability to generate samples beyond the support of the training data, particularly in out-of-distribution (OOD) regions, highlighting its strong generalization capability.
Figure 7 presents the results of the material generation task conditioned on a single property. Panels (a) and (e) show the density distributions of the target properties for the stable samples generated by UniGenX, alongside the corresponding label distributions in the fine-tuning dataset for magnetic (a) and mechanical (e) properties. The target value specified for UniGenX is indicated by a black dashed line. Panels (b–d) and (f–h) visualize representative S.U.N. structures generated under magnetic (b–d) and mechanical (f–h) property conditions, each annotated with the reduced formula, space group, and the predicted property value. Notably, UniGenX demonstrates the ability to generate samples beyond the support of the training data, particularly in out-of-distribution (OOD) regions. This highlights the model’s strong generalization capability—its ability to push forward from one distribution to another.
During experiments, we observed that the temperature and top-p value of UniGenX have a significant impact on the quality of material generation. Higher temperature and top-p values lead to increased diversity among generated samples and a greater proportion of novel materials. In contrast, lower values result in property distributions that are more sharply concentrated around the target. Under the optimal setting, UniGenX achieves an inference speed of 0.315 seconds per S.U.N. structure, demonstrating its practical efficiency for structural generation.
Figure 8: Generating materials with multi-objective constraints. This figure illustrates the performance of UniGenX on the task of material generation under multiple property constraints. In panel (a), gray dots represent the distribution of labeled training data, while green dots denote the distribution of samples generated by our model. The yellow-shaded region indicates the domain defined by the target constraints for this task ($³2.3\,\text{g/cm\textthreesuperior}\leq\text{density}\leq 2.7\,\text{g/cm\textthreesuperior}$, $\text{heat capacity}>1\,\text{J/g\textperiodcentered K}$, $E_{\text{Hill}}>200\,\text{GPa}$). Panel (b) provides a detailed view of how the generated samples populate the constrained property space. Panel (c) presents selected examples of SUN (Stable, Unique, and Novel) structures among the generated samples. Out of 1,262 generated candidates, our model successfully identified 436 stable structures that satisfy all specified property constraints. Four of these were confirmed by DFT to be novel, thermodynamically stable compounds not present in existing databases.
Figure 8 illustrates the performance of UniGenX on the task of multi-property-constrained material generation. In panel (a), gray dots represent the distribution of labeled training data, while green dots denote the distribution of structures generated by our model. The yellow-shaded region defines the desired target domain: $³2.3,\text{g/cm\textthreesuperior}\leq\text{density}\leq 2.7,\text{g/cm\textthreesuperior}$, $\text{heat capacity}>1,\text{J/g\textperiodcentered K}$, and $E_{\text{Hill}}>200,\text{GPa}$. Panel (b) zooms in on the coverage of the generated samples within this constrained property space, and panel (c) showcases selected SUN structures among the successful generations. Out of $1262$ generated candidates, our model identified $436$ structures predicted to be both stable and compliant with all target property constraints.
To further evaluate novelty, we conducted a rapid screening procedure based solely on elemental composition, filtering out structures present in the training set. This process yielded 11 novel candidates, of which 4 were confirmed through DFT calculations to satisfy all property requirements while also exhibiting thermodynamic stability. None of these four materials were found in the Materials Project database, suggesting they may represent previously unreported compounds.
Although this screening was intentionally limited in scope for computational efficiency, it demonstrates the model’s ability to generate valid, constraint-satisfying, and potentially novel materials. A more exhaustive comparison across the full set of generated structures would likely uncover additional unique candidates, offering further evidence of UniGenX’s potential in accelerating inverse material design under multi-objective constraints.